OpenStudy (anonymous):

i need help please. Using complete sentences, explain how to find the average rate of change for f(x) from x=4 to x=7 standard for is f(x)=2x^2-32x+64 vertex is y=(x^2-16x+32)

OpenStudy (tkhunny):

Needs work, but it's a start... What happenin' to the function, hey, from four to seven. Dude, we ain't gotta' go all the way up to Heaven. Where the function be when my Homes is four? Just write the answer down. Maybe on the floor. Make the next stop all the way to seven. That's where we gotta stop 'cause that is given. What's the difference, man, we must get on that subtraction. We almost done, I'm seeing the end o' this action. What's the average rate? Oh, we call out division. Do it right the first time. No time for revision. The function changes WHAT? We subtract for that! My Man move three. Don't keep under your hat. Divide the change of function by that nice little three. Show it to the teacher and then you're free. Complete enough sentences?

OpenStudy (anonymous):

what im sorry im so lost and confuse?

OpenStudy (tkhunny):

Those are complete sentences. That's the description of how to find the average rate of change.

OpenStudy (anonymous):

can you show me

OpenStudy (anonymous):

i forgot to give you this info standard for is f(x)=2x^2-32x+64 vertex is y=(x^2-16x+32)

OpenStudy (tkhunny):

I already showed you the complete sentences. Kids these days! \(\dfrac{f(7) - f(4)}{7-4}\) -- That's it.

OpenStudy (anonymous):

lmao i am so sorry im still lost math just aint my thing

OpenStudy (tkhunny):

...or RAP, apparently. What does f(7) mean?

OpenStudy (anonymous):

i dont know

OpenStudy (tkhunny):

Then you cannot solve this problem. Oh well, at least you have my complete sentences. I dare you to show them to your teacher.

OpenStudy (anonymous):

well thanxs for helping

OpenStudy (tkhunny):

You have a value for x = a. If you have a function, f(x), the f(a) means you evaluate the function at x = a. It's just a substitution problem. Why have you been given this problem if you have not also been given the tools to solve it? Is this a placement exam? If so, you should miss this problem so that you can be assigned to the correct course where you will get the right exposure.